40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3942     Entry time: Wed Nov 17 23:45:20 2010
Author: Jenne 
Type: Update 
Category: SUS 
Subject: A bad day for suspensions 

[Jenne, Suresh]

Today has been a downright miserable day in the world of suspension work. Thumbs down to that: 

Yesterday, we had glued 2 full sets of magnets to dumbbells.  Today, half of those broke.  I think I put too thin of a layer of glue on the magnets when gluing them to the dumbbells.  All magnet/dumbbell assemblies should pass the test of being picked up by the dumbbell while the magnet is stuck to the optical table or a razor blade.  6 of the 12 magnets failed this test. Suresh soaked the dumbbells that had been used in acetone, and scrubbed them, so we can reuse them when we reglue things tomorrow.  By some miracle, we have exactly one full set intact (for each set of 6, we need 4 of one direction and 2 of the other).  This was frustrating, but not yet a deal-breaker.  That part comes next....

I got ETMU05 nicely aligned in the magnet gluing fixture, and then was on my last check of whether the side magnets would be glued in the correct place when I realized that the fixture is all wrong for the ETMs.  This final check was added to the procedure after the drama with the ITMs of having the side magnets glued incorrectly as a result of the fixture being specific to the wedge angle of the optic.  Kiwamu and I had set the fixture to be just right for the ~1deg wedge corner station optics, but the ETMs have a 2.35deg wedge (according to the Coastline spec sheet, which is consistent with our measurements when placing the guiderod and standoffs).  Suresh and I need to reset the height of the optic in the fixture using more teflon sheets, but we don't have a whole lot of options ready in the cleanroom.  We're going to cut some more pieces and ask Bob to clean them tomorrow.  Since the way the fixture holds the teflon is a little hoaky, Suresh suggested just resting the optic on teflon pads, rather than screwing the teflon to the fixture, and then putting the optic on the pads.  We'll try Suresh's method tomorrow, and hopefully it will be pretty easy. 

At least the guiderods and standoffs were successfully glued to the optics....

Here's the updated Status Table.  I don't think we're going to be able to have an ETM ready for the chambers early next week, but we should still be able to have both ready for the Monday after Thanksgiving.  The spring plungers arrived today, and were given immediately to Bob and Daphen for cleaning.

StatusTable.png

ELOG V3.1.3-